The net worth of senior citizens is normally distributed with mean equal to $225,000 and standard deviation equal to $35,000. What is the probability that a single randomly selected senior citizen will have a net worth less than $300,000
Answer by stanbon(75887)

The net worth of senior citizens is normally distributed with mean equal to $225,000 and standard deviation equal to $35,000. What is the probability that a single randomly selected senior citizen will have a net worth less than $300,000
-------
z(300,000) = (300,000-225000)/35000 = 75/35 = 2.1429
---
P(x > 300,000) = P(z > 2.1429) = 0.0161
============
Cheers,
Stan H.